# Cross-Resolution SAR Target Detection Using Structural Hierarchy Adaptation and Reliable Adjacency Alignment

### v1

@@ V-F2 Discussions on Hyperparameter of RSAA
Fig. 17 illustrates the impact of the hyperparameter $r$ on detection performance. In the Aircraft LR→HR task, as $r$ increases from 0 to 5, both F1 and mAP rise to their highest values. However, when $r$ continues to increase beyond this point, neither F1 nor mAP improves further, suggesting that excessive number of source domain neighbors offer limited benefits for the discriminative performance improvement. For the Vehicle LR→HR task, the optimal F1 and mAP are observed when $r=3$, with a slight decrease in performance as $r$ increases further. These results demonstrate that, during the adjacency semantic alignment process, selecting a moderate number of source domain neighbors can effectively balance the strength of alignment and the influence of noise, thereby maximizing detection performance. In contrast, using too few or too many source neighbors does not provide additional performance improvements.
@@ VI Conclusion
In this paper, we present a novel SAR target detection framework (CR-Net) that effectively addresses the challenges of cross-resolution SAR detection. CR-Net achieves reliable cross-resolution detection by incorporating evidential learning theory and target scattering structure correlations, thereby evolving into two key modules: Structure-induced Hierarchical Feature Adaptation (SHFA) and Reliable Structural Adjacency Alignment (RSAA). Specifically, By exploring multi-mode structures within the domain, CR-Net integrates SHFA to perform fine-grained distribution adaptation under the guidance of structure similarity, thereby alleviating feature misalignment and enhancing the interpretability of the adaptation process. Furthermore, RSAA is integrated to improve the discriminability and generalization ability of CR-Net by transferring the discriminative information from source reliable adjacency samples. This dual approach enables CR-Net to significantly promote positive transfer while avoiding negative transfer. Extensive experimental results demonstrate that CR-Net substantially improves the performance of cross-resolution SAR target detection. In the future, we will continue to explore how to leverage foundation models to further enhance cross-resolution target detection performance, thereby promoting the practical application of this technology.

### v2

@@ V-F2 Discussions on Hyperparameter of RSAA
Fig. 18 illustrates the impact of the hyperparameter $r$ on detection performance. While a larger $r$ enables improved semantic alignment by capturing richer neighbor semantic information, it also increases the risk of introducing noise-instances and error accumulation. For the Aircraft LR→HR task, both F1 and mAP steadily improve as $r$ increases from 0 to 5, reaching their highest values at $r=5$. In contrast, for the Vehicle LR→HR task, optimal performance is achieved at $r=3$. Notably, further increasing $r$ beyond these points leads to performance declines in both datasets. This phenomenon occurs because an excessively large $r$ may introduce semantically inconsistent noise-instances, leading to error accumulation during semantic alignment. Moreover, the optimal $r$ for the Vehicle dataset is lower than that for the Aircraft dataset. The difference can be attributed to the fact that vehicle targets typically appear in urban scenes with more complex and diverse contexts, which increases the proportion of structurally similar but semantically different noise-instances (e.g., containers, air conditioner units). Based on the above analysis, the optimal value of $r$ varies across different datasets and should be carefully selected to maximize detection performance. In this paper, we set $r=5$ for the Aircraft dataset and $r=3$ for the Vehicle dataset by default.
@@ VI Conclusion
In this paper, we present a novel SAR target detection framework (CR-Net) that effectively addresses the challenges of cross-resolution SAR detection. CR-Net achieves reliable cross-resolution detection by incorporating evidential learning theory and target scattering structure correlations, thereby evolving into two key modules: Structure-induced Hierarchical Feature Adaptation (SHFA) and Reliable Structural Adjacency Alignment (RSAA). Specifically, By exploring multi-mode structures within the domain, CR-Net integrates SHFA to perform fine-grained distribution adaptation under the guidance of structure similarity, thereby alleviating feature misalignment and enhancing the interpretability of the adaptation process. Furthermore, RSAA is integrated to improve the discriminability and generalization ability of CR-Net by transferring the discriminative semantics from source reliable adjacency samples. This dual approach enables CR-Net to significantly promote positive transfer while avoiding negative transfer. Extensive experimental results demonstrate that CR-Net improves the performance of cross-resolution SAR target detection. In the future, we will continue to explore how to leverage foundation models to further enhance cross-resolution target detection performance, thereby promoting the practical application of this technology.